# Learning Efficient Representations of Neutrino Telescope Events

### v2

@@ 5.3 Runtime Efficiency
In terms of forward pass encoding runtime, om2vec proves to be an efficient method, particularly when compared to classical approaches like the AGMM. Table 2 presents the average per-PATD runtime for various encoding methods alongside their latent dimensionalities. om2vec also significantly benefits from GPU acceleration, unlike the AGMM. Even when running on a CPU, om2vec is an order of magnitude faster. This is particularly important when running in resource-constrained environments such as IceCube’s lab at the South Pole. With GPU acceleration, this speed advantage increases to two orders of magnitude. It is also worth noting that PATDs can be batched and processed in parallel on the GPU, which would further speedup the average runtime. Additionally, om2vec scales well with increasing latent dimensionality, allowing users to expand the latent space with minimal runtime impact.
We also observe runtime improvements when running angular reconstruction algorithms on events using om2vec-represented PATDs instead of full-timing information. SSCNN (Full) took approximately 8.5s to process 20,000 events. In contrast, SSCNN (om2vec) took only 2.1s for the same number of events, representing a significant speedup. All GPU runtime tests were conducted on an NVIDIA A100 80GB.
@@ 6 Conclusions and Future Directions
This work introduces om2vec, a novel approach utilizing transformer-based VAEs to efficiently represent neutrino telescope events. Our method addresses the significant challenges posed by the large size, high dimensionality, and variance of data in neutrino telescopes. By learning compact and descriptive latent representations, om2vec offers several advantages for downstream analysis tasks.
@@
Availability: Source code, datasets, and pre-trained checkpoints for om2vec on Prometheus events is made available on GitHub. Additionally, as a future direction, om2vec is will be integrated into GraphNeT [28], an open-source deep learning pipeline widely utilized by various collaborations in the neutrino telescope community.

### v3

@@ 5.3 Runtime Efficiency
In terms of forward pass encoding runtime, om2vec proves to be an efficient method, particularly when compared to classical approaches like the AGMM. Table˜2 presents the average per-PATD runtime for various encoding methods alongside their latent dimensionalities. om2vec also significantly benefits from GPU acceleration, unlike the AGMM. Even when running on a CPU, om2vec is an order of magnitude faster. This is particularly important when running in resource-constrained environments such as IceCube’s lab at the South Pole. With GPU acceleration, this speed advantage increases to two orders of magnitude. It is also worth noting that PATDs can be batched and processed in parallel on the GPU, which would further speedup the average runtime. Additionally, om2vec scales well with increasing latent dimensionality, allowing users to expand the latent space with minimal runtime impact.
The use of latent representations also leads to notable runtime improvements during downstream reconstruction, such as angular reconstruction, compared to models that rely on full-timing information. SSCNN (Full) took approximately 8.5s to process 20,000 events. In contrast, SSCNN (om2vec) took only 2.1s for the same number of events, representing a significant speedup. All GPU runtime tests were conducted on an NVIDIA A100 80GB.
@@ 6 Conclusions and Future Directions
This work introduces om2vec, a novel approach utilizing transformer-based VAEs to efficiently represent the PATDs of neutrino telescope events. Our method addresses the significant challenges posed by the large size, high dimensionality, and variance of data in neutrino telescopes. By learning compact and descriptive latent representations, om2vec offers several advantages for downstream analysis tasks.
@@
Availability: Source code, datasets, and pre-trained checkpoints for om2vec on Prometheus events is made available on GitHub. Additionally, as a future direction, om2vec is will be integrated into GraphNeT [Søgaard2023], an open-source deep learning pipeline widely utilized by various collaborations in the neutrino telescope community.